OpenStudy (anonymous):

use the given trigonometric identity to set up a u-substitution and then evaluate the indefinite integral: (intergal of) 4cos^2x dx, with the identity cos2x=1-2cos^2x

hartnn (hartnn):

@mulline hi :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) from cos2x=1-2cos^2x try to isolate cos^2 x.

OpenStudy (anonymous):

Thanks I totally get it now!

hartnn (hartnn):

ohh, good :)

OpenStudy (anonymous):

can you help me with this one

hartnn (hartnn):

yeah, you can use, \(a^2-b^2 = (a+b)(a-b)\) here, a = cos^2x , b = sin^2 x

hartnn (hartnn):

and use the fact that : \(\sin^2x+\cos^2x=1\)

OpenStudy (anonymous):

youre a genius.

hartnn (hartnn):

not actually :P math identities are genius :)